Social science. In a town election, the winning candidate had 220 more votes
than the loser. If 810 votes were cast in all, how many votes did each candidate
receive?
Iam so bad with word problems please help.
Debbie

Answer by scott8148(6628)

x=loser, x+220=winner ... x+x+220=810 ... 2x=590 ... x=295 ... x+220=515